# The Full-scale Assembly Simulation Testbed (FAST) Dataset

### v1

@@ Introduction
Using FAST, we have conducted a within-subject study with 108 participants (50 females, 56 males, 2 non-binary) that involved learning how to assemble two distinct full-scale structures (see Fig. The Full-scale Assembly Simulation Testbed (FAST) Dataset). The experimental procedure involved completing an online screener that included a demographics survey, completing an in-VR tutorial for how to use FAST, using FAST to learn how to construct one of the two structures, constructing the structure in real life from memory, using FAST to learn how to construct the other structure, constructing it in real life, and an exit survey. Data collected from this study include the tracking and interaction data from FAST, participant demographics, and multiple questionnaires, including the System Usability Scale (SUS) [1], Simulator Sickness Questionnaire (SSQ) [12], NASA Task Load Index (TLX) [5], and the Spatial Presence Experience Scale [6].
@@ 2.1 Procedure
In addition to the physical tasks asked of our participants, they were given some common questionnaires throughout the study including the System Usability Scale (SUS) [1], Simulator Sickness Questionnaire (SSQ) [12], NASA Task Load Index (TLX) [5], and the Spatial Presence Experience Scale [hartmann2015spatial]. The ordering of the administration of these questionnaires was designed to ensure that we would acquire immediate before and after measures of SSQ, and per-build measures of TLX, SUS, and SPES.
@@ 2.4 Collected Data
FAST collected two sets of data for each participant, one dataset per assembly task (i.e., A and B). Each set included data about the local tracking position and rotation of the VR tracked devices every frame at 90Hz for the duration of the assembly task. We have open-sourced the collective dataset on GitHub for other researchers to use [^3] https://github.com/xrtlab/FAST-Dataset.

### v2

@@ Introduction
Using FAST, we have conducted a within-subject study with 108 participants (50 females, 56 males, 2 non-binary) that involved learning how to assemble two distinct full-scale structures (see Fig. The Full-scale Assembly Simulation Testbed (FAST) Dataset). The experimental procedure involved completing an online screener that included a demographics survey, completing an in-VR tutorial for how to use FAST, using FAST to learn how to construct one of the two structures, constructing the structure in real life from memory, using FAST to learn how to construct the other structure, constructing it in real life, and an exit survey. Data collected from this study include the tracking and interaction data from FAST, scores quantifying participants’ real-world performance, participant demographics, and multiple questionnaires, including the System Usability Scale (SUS) [1], Simulator Sickness Questionnaire (SSQ) [12], NASA Task Load Index (TLX) [5], and the Spatial Presence Experience Scale [6].
@@ 2.1 Procedure
In addition to the physical tasks asked of our participants, they were given some common questionnaires throughout the study including the System Usability Scale (SUS) [1], Simulator Sickness Questionnaire (SSQ) [12], NASA Task Load Index (TLX) [5], and the Spatial Presence Experience Scale [6]. The ordering of the administration of these questionnaires was designed to ensure that we would acquire immediate before and after measures of SSQ, and per-build measures of TLX, SUS, and SPES.
@@ 2.4 Collected Data
FAST collected two sets of data for each participant, one dataset per assembly task (i.e., A and B). Each set included data about the local tracking position and rotation of the VR tracked devices every frame at 90Hz for the duration of the assembly task. We have open-sourced the collective dataset on Hugging Face for other researchers to use [^3] https://huggingface.co/datasets/XraiLab/FAST-Dataset.
The photographs of the participant’s recreations were used to model them as 2D graphs, with pipes serving as the edges of the graph and connectors serving as the nodes. The ground-truth assembly for each task was also modeled as a graph. A modified A* search algorithm [25] was then used to identify the most efficient set of operations to transform the graphs of the participants’ recreations into the graph of the ground-truth assembly for the corresponding tasks.
Before employing the algorithm, an optimization step was used to account for ’free moves’, or actions that reoriented the graph but did not incur a ’cost’ towards the overall solution. ’Flipping’, or reversing the connections of a piece, and ’swapping’, exchanging the positions of two pipes or connectors of the same type (color and shape respectively), were the available free actions. An exhaustive, parallel search generated every possible combination of free moves for the graph. The combination of free moves that resulted in the smallest possible Manhattan distance between the adjacency matrices of the recreation and the ground-truth graph was selected. The assembly was then compared with ground-truth, to identify those pipes and connectors attached to the correct pieces, but rotated incorrectly. As the substructure these pieces formed would not be modified by attachments or detachments, a flat cost of 2 was assigned for each rotation required. The value incurred from this step served as the initial start cost of the overall solution for the A* algorithm.
The ’costly’ moves allowed by the algorithm included only attachments and detachments, each incurring a cost of 1 per move. The heuristic function guiding the search was the Manhattan distance between the adjacency matrix of the recreation, with the chosen operations applied, and the adjacency matrix of the ground truth assembly. A final penalty of 1 per unused piece was applied. The final number of operations required for each participant’s recreation was recorded as their score for the corresponding task.